OpenStudy (anonymous):

difference of two geometric series. infinity(summation symbol) n=1 [(0.8)^(n-1) - (0.3)^n]

OpenStudy (aravindg):

can u pls se eqn editor?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(0.8)^{n-1}-(0.3)^n\]

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} [(0.8)^{n-1} -(0.3)^{n}]\]

OpenStudy (aravindg):

i suggest u open the brackets and evaluate sigma separatel and then add

OpenStudy (anonymous):

I get (1-0.3)+ (0.8-0.9) Which is wrong, obviously.So do I find first find what both series converge to then subtract them?

OpenStudy (anonymous):

Never mind. I think I got it

OpenStudy (anonymous):

Still struggling...I guess the 1-.3 is corrected but I'm not quite sure what to put in the denominator....please help

OpenStudy (anonymous):

I have the solution to the question. it's \[1/(1-0.8) -0.3/(1-0.3) = 5- 3/7 = 32/7 \] BUT WHY?

OpenStudy (aravindg):

wait a sec ..lemme do this on paper then i will post my answer

OpenStudy (anonymous):

Thanks

OpenStudy (aravindg):

yep !!

OpenStudy (aravindg):

:(

OpenStudy (aravindg):

I AM TERRIBLY SAD !!

OpenStudy (aravindg):

I WROTE A BIG ANSWER AND SUDDENLY GOOGLE CHROME SHOWS "AW SNAP "RELOAD THE PAGE :(

OpenStudy (anonymous):

Ohhh I'm so sorry....darn it Google Chrome!!!!

OpenStudy (aravindg):

i typed it so nicely for u :( .....anyway i will try to draw the answer i cant bare another latex writing

OpenStudy (anonymous):

sure, anything will help. Thanks for the effort

OpenStudy (aravindg):